OpenStudy (anonymous):

PLEEEEEASE help and explain! :( Vladimir is making tapioca pudding. He started 2 3/4 with fluid ounces of milk. As he cooked, he used a tablespoon of milk, which is 1/2 of a fluid ounce, x times. When he finished, he had at least 1 1/4 fluid ounces of milk left. Which phrase best describes the number of times Vladimir used a tablespoon of milk?

jimthompson5910 (jim_thompson5910):

please post the full problem

jimthompson5910 (jim_thompson5910):

thank you

OpenStudy (anonymous):

sorry lol

jimthompson5910 (jim_thompson5910):

something seems missing in this sentence ` When he finished, he had at least fluid ounces of milk left`

jimthompson5910 (jim_thompson5910):

I'm guessing a number or algebraic expression is in there (or supposed to be)?

OpenStudy (anonymous):

fixed it

jimthompson5910 (jim_thompson5910):

oh ok, 1 and a 1/4, thanks

jimthompson5910 (jim_thompson5910):

when they state `Which phrase best describes the number of times Vladimir used a tablespoon of milk?` I'm assuming that means you're given a list of choices?

OpenStudy (anonymous):

the options are : at least 3 times at most 3 times at most 8 times at least 1 1/2 times

jimthompson5910 (jim_thompson5910):

ok let me think

jimthompson5910 (jim_thompson5910):

do you agree that 2&3/4 is equivalent to 2.75 in decimal form?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

1/2 = 0.5 and 1&1/4 = 1.25

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so we start off with 2.75 fl oz of milk we subtract off 0.5 fl oz per tablespoon, there are x of these tablespoons, so 0.5x is subtracted off

jimthompson5910 (jim_thompson5910):

2.75 - 0.5x is the amount left over after using x tablespoons

jimthompson5910 (jim_thompson5910):

we want this to be at least 1.25 fl oz, so that means we have this inequality \[\Large 2.75 - 0.5x \ge 1.25\] do you see how to isolate x?

OpenStudy (anonymous):

\[x \le 3\]

OpenStudy (anonymous):

is that correct?

jimthompson5910 (jim_thompson5910):

very good, so that means x is at most 3 therefore, he must use at most 3 tablespoons (3 is the upper limit)

OpenStudy (anonymous):

okay thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome